OpenStudy (whynot):

{(-1,9),(2,9),(5,135)} Write the equation in standard form y=ax^2+bx+c

OpenStudy (anonymous):

So we have to fit a quadratic function?

OpenStudy (whynot):

yes

OpenStudy (saifoo.khan):

Hey Lahore. long time no C.

myininaya (myininaya):

so we have the following: \[f(-1)=9 ; f(2)=9 ; f(5)=135\] we need to do some plugin'

myininaya (myininaya):

\[f(-1)=a-b+c=9\] \[f(2)=4a+2b+c=9\] \[f(5)=25a+5b+c=135\]

OpenStudy (anonymous):

solve using gaussian elimination!!

OpenStudy (anonymous):

1 0 0 7 0 1 0 -7 0 0 1 -5

OpenStudy (anonymous):

a = 7, b = -7, c = 5. or am I mistaken I will draw the graph

OpenStudy (anonymous):

It's wrong :(

OpenStudy (whynot):

its alright @agd :) i kinda get it as myininaya is explaining

OpenStudy (anonymous):

I will call numpy's quadratic fitting function and find the answer

OpenStudy (anonymous):

my numpy implementation is broken (I'/m using a 64 bit platform so it won't work) I will code gradient descent myself!!!

OpenStudy (whynot):

thank you agdgdgdgwngo :) that is helpful .extra medal for you

OpenStudy (anonymous):

I've found an answer. a = 3.75 b = 0 c = 5.25

OpenStudy (anonymous):

It doesn't fit 135 :(

myininaya (myininaya):

\[a-b+c=9\] \[4a+2b+c=9\] \[25a+5b+c=135\] multiply equation one by 2 and add to equation two giving us the new set of equations \[a-b+c=9\] \[6a+0+3c=27\] \[25a+5b+c=135\] multiply equation one by 5 and add to equation three giving us the new set of equations \[a-b+c=9\] \[6a+0+3c=27\] \[30a+0+6c=180\] lets simplify little before doing anything else \[a-b+c=9\] \[2a+0+c=9\] \[5a+0+c=30\] Ok multiply equation three by -1 and add to equation one and two to get the new set of equations \[-4a-b+0=-21\] \[-3a+0+0=-21\] \[5a+0+c=30\] now simplify again \[-4a-b+0=-21\] \[a+0+0=7\] \[5a+0+c=30\] Ok multiply second equation by 4 and add to equation one \[0-b+0=7\] \[a+0+0=7\] \[5a+0+c=30\] Ok multiply second equation by -5 and add to equation three \[0-b+0=7\] \[a+0+0=7\] \[0+0+c=-5\] now lets simplify alittle \[0+b+0=-7\] \[a+0+0=7\] \[0+0+c=-5\]

OpenStudy (anonymous):

wow a lot of work there myininaya :(

myininaya (myininaya):

lol

myininaya (myininaya):

a=7; b=-7; c=-5

OpenStudy (anonymous):

those don't work either :(

myininaya (myininaya):

i don't believe you

OpenStudy (anonymous):

at least mine fits (1, 9) and (-1, 9)

myininaya (myininaya):

let me check lol

OpenStudy (whynot):

wow myininaya that was awesome!!!

OpenStudy (anonymous):

she used gauss-jordan elimination :-D

OpenStudy (whynot):

:D happy! thankyou so much myininaya and agdgdgdgwngo

OpenStudy (anonymous):

wait myininaya is right!!

OpenStudy (anonymous):

I wrote 1 instead of 2 :-P

myininaya (myininaya):

i checked it it works lol

myininaya (myininaya):

but i guess i'm late in the checking part

OpenStudy (anonymous):

you could also just do \[\Theta = (X^TX)^{-1}*X^{T} Y\]

myininaya (myininaya):

Show off! I think this is just a algebra question?

OpenStudy (whynot):

man you guys are so smart:) keep it up

myininaya (myininaya):

so agdgd are you going to know me what pseudoinverse means?

OpenStudy (anonymous):

sure

myininaya (myininaya):

lets go to your question